OpenStudy (elisaneedshelp):

What is the value of 8 4/3? Please dont think Im lazy. I actually dont understand. Please. Medal and Fan

OpenStudy (chaise):

\[\ 8 \frac{ 4 }{ 3 }\] Are you asking how we can simplify this?

OpenStudy (elisaneedshelp):

Maybe, i really dont get what its asking but i can provide you with the answer choices A. 2 B. 6 C. 8 D. 16

OpenStudy (chaise):

Am not sure, none of your answers match up with what I think would be the correct answer, which is 9.333 or 28/3

OpenStudy (amistre64):

maybe 8/1 --- 4/3

OpenStudy (amistre64):

you need to define what the operation between the terms is, otherwise this is just guesswork

OpenStudy (amistre64):

if: \[8\div \frac43\] \[\frac84\div \frac13\] \[2\div \frac13\] etc...

OpenStudy (amistre64):

cross 'division' if there was ever such a name lol

OpenStudy (elisaneedshelp):

oh my God x_x

OpenStudy (elisaneedshelp):

If the fraction looks like an exponent then would we be talking about the same thing?

OpenStudy (amistre64):

no, exponents are not division ...

OpenStudy (amistre64):

\[8^{4/3}=(8^4)^{1/3}\]

OpenStudy (elisaneedshelp):

So then what do i do?

OpenStudy (amistre64):

memorize some rules and learn to prime factor .....

OpenStudy (amistre64):

we could solve it simply by applying:\[a^{pq}=(a^p)^q\] and knownig the prime factors of 8

OpenStudy (elisaneedshelp):

Thank you I have to go now. Ill try to understand more when i get back to this.

OpenStudy (amistre64):

good luck